Solve
4x^3+x^2-36x-9=0
Answers
A) 3,(1/4)
B) -9,9,-(1/4)
C) -3,3,-4
D) -3,3,-(1/4)
ps the ^ means to the power of....thanks soo much

Answer by edjones(8007)

4x^3+x^2-36x-9=0
x^2(4x+1)-9(4x+1)=0 Factoring by grouping.
(x^2-9)(4x+1)=0
(x+3)(x-3)(4x+1)=0
x=3, x=-3, x=-1/4
.
Ed